OpenStudy (anonymous):

Help!! let h(x) = (g o f)(x) = 1 / (x+3)^2 Which of the following could be a possible decomposition of h(x)?

OpenStudy (phi):

how about g= 1/x^2 and f= (x+3) ?

OpenStudy (anonymous):

are you sure it's not the other way around? I got that same answer but then i had doubts haha @phi

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

why dont i see the answer choices O_O